Hi
John had 100 dollars to spend on cakes. The shop had an ad with the following prices.
3 dollars a cake
Buy 5 cakes get 2 for free.
What is most number of cakes that he can buy with all his money.
Thanks

Found 2 solutions by Theo, Edwin McCravy:
Answer by Theo(13342)

the cost per cake = 3 dollars.
he has 100 dollars to spend.
100 / 3 = 33 cakes that he can buy with his 100 dollars.
he would pay a total of 3 * 33 = 99 dollars and have 1 dollar left over.
33 cakes divided by 5 = 6 sets of 5 cakes with 3 cakes left over.
6 * 5 = 30 + 3 = 33 total cakes.
for every set of 5 cakes, he gets 2 for free.
that means he gets 6 * 2 = 12 cakes for free on top of the 33 that he bought.
he will be able to gt 33 + 12 = 45 cakes for his money.
the 1 dollar he has left over can't be used to buy any more cakes, since the minimum cost per cake is 3 dollars and it is assumed they only sell whole cakes, not parts of a cake.

for his 100 dollars, he can get a maximum of 45 cakes, 33 that he paid for at 3 dollars each and 12 that he got for free.








Answer by Edwin McCravy(20059)

John had 100 dollars to spend on cakes. The shop had an ad with the following
prices.

3 dollars a cake

Buy 5 cakes get 2 for free.

What is most number of  cakes that he can buy with all his money.
First of all, this is the old "buy some and get some free" trick which has been
played on customers for many years.  [The one I see most often is "Buy 3 tires
at the regular price and get the 4th tire free".  

So let's eliminate the trick played.  If he buys 5 cakes for $15 and get 2 free,
the bottom line is that he's paid $15 and gotten 7 cakes.  So the merchant
might as well have said "$3 a cake, or 7 cakes for $15".

So we divide $15 into $100 to see how many groups of 7 cakes he could get:
                                     
                          6 
                     15)100
                         90
                         10

That would give him 6∙7 = 42 cakes.  That remainder of 10 tells us that he would
have $10 remaining.  That's not enough money remaining to buy 5 more cakes at $3
a cake, so he could not get any free cakes.  But he could still use it to get a
few more cakes even if he couldn't get any free ones.

                        3
                     3)10
                        9
                        1

So he could get 3 more cakes, or 42+3=45 cakes and have 1 dollar remaining. 

Edwin